If you are rolling a die and counting aces (an ace on a die is a 1), how many times should you roll the die in order to have a better than 80% chance of getting at least one ace?
Answer by richwmiller(17219)

80%=4/5
4/5=x/6
24=5x
24/5=x
4.8=x
so you need to roll the die at least 5 times for a better than 80% chance of rolling an ace.
each roll is 1/6 chance
x/6>4/5